# Long-range interacting systems are locally non-interacting

### v1

@@ Abstract
Long-range interacting systems are frequently considered in the context of spins, which are arranged on a $d$-dimensional lattice [15, 16, 17], and which interact pairwise with couplings following a power-law decay [cf. Fig. 1(a)]. Exploring these systems is challenging (see also results in classical settings [18, 19, 20, 21]), as state-of-the-art numerical methods face difficulties implementing long-range interactions and capturing correlations [22, 23, 24]. Moreover, only a few analytical results have been established beyond the case of quadratic theories [12] or systems with infinite-range interactions [25, 26, 27], let alone for nonequilibrium open quantum dynamics [28, 29, 30, 31, 32, 33, 34, 35].
In this paper, we rigorously derive the emergent dynamics of long-range interacting open quantum systems, in the strong long-range regime [1]. We consider generic many-body systems, in arbitrary dimension, with power-law decaying interactions characterized by an exponent $\alpha\leq d$ and a dynamical generator that is extensive in the particle number [36, 12, 37]. We demonstrate that, in the thermodynamic limit, the long-range dynamics is exactly captured by an effective nonlinear mean-field theory [cf. Fig. 1(b)].
@@
The superscripts highlight the sites on which the operators act. The first sum contains (possibly inhomogeneous) single-site terms $h_{k}$, while the second one describes long-range interactions, encoded in the (possibly inhomogeneous) two-body operators $V_{jk}$. The interaction strength decays algebraically with the distance, through the interaction-range exponent $\alpha$ [cf. Fig. 1(a)]. For $\alpha\leq d$, the double sum in Eq. (1) becomes super-extensive in $N$, leading to unphysical behavior in the thermodynamic limit $N\to\infty$ [26]. To prevent this, we include the so-called Kac factor, $c_{\alpha}^{N}=\sum_{j\in\Gamma}\|\textbf{r}_{j}\|_{\Gamma}^{-\alpha}$ [see Eq. (1)], which restores the extensivity of the Hamiltonian [36]. For large $N$, the Kac factor scales as [26]
@@
$\gamma_{jk}^{\mu}=\kappa_{j}^{\mu}\delta_{jk}+\frac{w_{jk}^{\mu}}{2c_{\alpha}^{N}[\delta_{jk}+(1-\delta_{jk})\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.$ (7)
@@
$\mathcal{L}_{N}=\sum_{k\in\Gamma}\mathcal{L}_{k}^{(k)}+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Gamma}\frac{\mathcal{L}_{jk}^{(jk)}}{[\delta_{jk}+(1-\delta_{jk})\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.$ (8)
The second term features an unconstrained double sum, so that the factor $1/2$ avoids double counting.
@@
$\begin{split}\dot{\varrho}_{N,t}^{(\Lambda)}&=\sum_{k\in\Lambda}\mathcal{L}_{k}^{(k)}[\varrho_{N,t}^{(\Lambda)}]+\frac{1}{c_{\alpha}^{N}}\sum_{j\in\Lambda,k\in\Lambda^{\prime}}\mathcal{C}_{jk}^{N}[{\varrho}_{N,t}^{(\Lambda\cup\{k\})}]\\ &+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}[{\varrho}_{N,t}^{(\Lambda)}]}{[\delta_{jk}/2+(1-\delta_{jk})\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.\end{split}$ (9)
@@
for $j\in\Lambda$, $k\in\Lambda^{\prime}$. The relation in Eq. (9) defines a hierarchy of differential equations, manifesting in ${\varrho}_{N,t}^{(\Lambda)}$ depending on the reduced states ${\varrho}_{N,t}^{(\Lambda\cup\{k\})}$ involving the particles in $\Lambda$ plus a particle in $\Lambda^{\prime}$ [25, 31, 45, 27].
@@
Our theorem shows that, for large $N$, reduced states of the many-body quantum system remain in product form, if they initially are, and are fully determined by the single-particle states $\tilde{\varrho}_{N,t}^{(j)}$. The latter evolve through a dissipative Hartree-like [25, 45] dynamics [cf. Eq. (13)]. Contrary to a standard (permutation-invariant) mean-field scenario, the generic situation considered so far does not allow for an immediate statement on the convergence, with $N$, of $\tilde{\varrho}_{N,t}^{(j)}$ to a well-defined limit $\tilde{\varrho}_{\infty,t}^{(j)}$. This is due to the allowed inhomogeneous parameters, or disorder, in the dynamical generator [cf. Eq. (1) and Eq. (7)]. Moreover, it is intrinsically related to the fact that long-range interactions are not permutation invariant. As a consequence, the convergence of reduced single-particle states is strictly tied to the convergence of the average over collisions in Eq. (13) (see Proposition 1 below), which is not obvious as in permutation-invariant cases. Even in the most general case, our theorem however permits the efficient analysis of the system as well as of convergence of the reduced single-particle states, e.g., via numerical integration of Eq. (13) for increasing $N$.
@@
We take a spin-$1/2$ chain ($d=1$) with open boundary conditions, as illustrated in Fig. 2(a), and unitary dynamics governed by the Hamiltonian in Eq. (1), with $h_{k}=0$ and $V_{12}^{(jk)}=\Omega\sigma^{(j)}_{x}\sigma^{(k)}_{x}$. The dissipative contribution is local, with
@@
with $\phi(\textbf{x})=\langle\sigma_{x}\rangle\Omega\theta(\alpha)\int_{0}^{1}{\rm d}\textbf{y}\,|\textbf{x}-\textbf{y}|^{-\alpha}$. Such an equation is benchmarked in Fig. 3(d).
The nonlinear and non-interacting character of the dynamics governed by Eq. (17) provides simple, yet rigorous, insights on the emergence of the long-lived quasi-stationary states observed in Refs. [26, 12]. There, a unitary quench from a state with $\langle\sigma_{x}\rangle=0$ was considered. Our results imply that, in the thermodynamic limit, the system must obey Eq. (17), with $\kappa=0$, so that an initial state with $\langle\sigma_{x}\rangle=0$ is a fixed point of the emergent dynamics. Therefore, for finite systems, the quantum many-body state must necessarily become longer and longer lived upon increasing the number of particles. [Fig. 3(b) displays dynamics since we consider a state with $\langle\sigma_{x}\rangle\neq 0$.]
Discussion.— Our theorem is very general and covers the cases of disordered single-site and interacting terms as well as of generic inhomogeneous dynamics. Importantly, it allows for the exploration of strong long-range interacting dynamics with a linear complexity in the system size [cf. Eq. (13)]. Our proposition demonstrates the impact of boundary conditions on the emergent dynamics. For periodic boundaries and translation-invariant dynamical generators, the system becomes permutation invariant in the thermodynamic limit. For open boundaries, it is instead described by an inhomogeneous mean-field theory (see, e.g., Refs. [49, 50, 51, 52] for inhomogeneous mean-field equations in related contexts). While Eqs. (13-15) may look abstract at first, they simply entail that the relevant equations of motion can be found by considering a product-state ansatz.
@@
Acknowledgments.— We are grateful to Alvise Bastianello for useful discussions. We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Unit FOR 5413/1, Grant No. 465199066, through the Research Unit FOR 5522/1, Grant No. 499180199. This project has also received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). FC is indebted to the Baden-Württemberg Stiftung for the financial support of this research project by the Eliteprogramme for Postdocs.
@@
Robert Mattes,1 Igor Lesanovsky,1,2 and Federico Carollo1
@@
$\mathcal{L}^{(\Lambda)}=\sum_{k\in\Lambda}\mathcal{L}^{(k)}_{k}+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}}{[\delta_{jk}+(1-\delta_{jk})\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,,$ (S3)
@@
The last inequality is obtained by shifting the origin of the lattice $\Gamma$ to $\textbf{r}_{j_{i}}$ and adding in each dimension auxiliary sites in such a way that all entries of $\textbf{r}_{k_{i}}$ range in $r_{k_{i}}^{\mu}\in[-L,L]$, not including $r_{k_{i}}^{\mu}=0$. This is the reason for the term $2^{d}$. Moreover, each sum over $j_{i}$ is bounded by $n+i-1$, where $n=|\Lambda|$ is the extension of the subset $\Lambda$. All these considerations give
@@
$\left\|e^{s\mathcal{L}^{(\Lambda)}}\circ\left[\mathcal{L}^{(\Lambda)}-\hat{\mathcal{L}}^{(\Lambda)}\right]\circ e^{(t-s)\hat{\mathcal{L}}^{(\Lambda)}}[Q^{(\Lambda)}]\right\|_{\rm tr}$ $\leq\left\|\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}\circ e^{(t-s)\hat{\mathcal{L}}^{(\Lambda)}}[Q^{(\Lambda)}]}{[\delta_{jk}+(1-\delta_{jk})\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\right\|_{\rm tr}\leq\left\|Q^{(\Lambda)}\right\|_{\rm tr}\frac{C|\Lambda|^{2}}{c_{\alpha}^{N}}\,.$
@@
$\|\mathcal{M}_{t}[\cdot]\|_{\mathrm{tr}}\leq\|\mathcal{L}_{1}\left[\cdot\right]\|_{\mathrm{tr}}+\int{\rm d}\textbf{z}\frac{1}{\|\textbf{x}-\textbf{z}\|_{\Gamma}^{\alpha}}\left\|\mathrm{tr}_{\textbf{z}}\circ\mathcal{L}_{12}\circ\left[\cdot\otimes\tilde{\varrho}_{\infty,t}^{(\textbf{z})}\right]\right\|_{\mathrm{tr}}\leq\left(D+\int{\rm d}\textbf{z}\frac{1}{\|\textbf{x}-\textbf{z}\|_{\Gamma}^{\alpha}}C\right)\|\cdot\|_{\mathrm{tr}}=M\|\cdot\|_{\mathrm{tr}}\,,$
@@
IV IV. Collision-model circuit and additional results for the long-range dissipative quench
@@
This formula shows that we can calculate the evolution of the reduced state $\varrho_{N,t}^{(\{j,k\})}$ as follows. First, we consider the interaction $\mathcal{U}_{jk}$ between the particles of interest. Then, we make both site $j$ and site $k$ sequentially interact with one particle in $\{j,k\}^{\prime}$ and tracing the latter out. After all particles in the remainder of the system have been considered, we can apply the dissipative contributions. This algorithm, schematically illustrated in Fig. (2)(b), allows us to investigate the many-body system with a complexity which is equivalent to that of a three-qubit system subject to order $N$ maps, which involve at most two qubits. In the numerical calculations, we have used the initial state $\varrho_{N}^{(j)}=\left(\begin{array}[]{cc}0.2&0.2-0.2i\\ 0.2+0.2i&0.8\end{array}\right)\,,$ for each site.
In Fig. S1(a,b), we show results for the quantum mutual information between the two sites, $\mathcal{I}(j:k)=\mathcal{S}(\varrho_{N,t}^{(\{j\})})+\mathcal{S}(\varrho_{N,t}^{(\{k\})})-\mathcal{S}(\varrho_{N,t}^{(\{j,k\})})$, with $\mathcal{S}(\varrho)=-\mathrm{tr}\left(\varrho\log(\varrho)\right)$ being the von Neumann entropy. We see that the quantum mutual information decreases with the system size in the strong long-range regime $\alpha\leq 1$. Its behavior as a function of $N$ [shown in Fig. S1(b)] is in principle not the same shown in the inset of Fig. 3(b) by the correlations $C_{zz}$, due to the definition of the quantum mutual information through entropy functions.

### v2

@@ Abstract
Long-range interacting systems are frequently considered in the context of spins, which are arranged on a $d$-dimensional lattice [15, 16, 17], and which interact pairwise with couplings following a power-law decay [cf. Fig. 1(a)]. Exploring these systems is challenging (see also results in classical settings [18, 19, 20, 21]), as state-of-the-art numerical methods face difficulties implementing long-range interactions and capturing correlations, in particular for open quantum systems and close to phase transitions [22, 23, 24]. Moreover, only a few analytical results have been established beyond the case of quadratic theories [12] or systems with infinite-range interactions [25, 26, 27], let alone for nonequilibrium open quantum dynamics [28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
In this paper, we rigorously derive the emergent dynamics of long-range interacting open quantum systems, in the strong long-range regime [1]. We consider generic many-body systems, in arbitrary dimension, with power-law decaying interactions characterized by an exponent $\alpha\leq d$ and a dynamical generator that is extensive in the particle number [38, 12, 39]. We demonstrate that, in the thermodynamic limit, the long-range dynamics is exactly captured by an effective nonlinear mean-field theory [cf. Fig. 1(b)]. Our study goes beyond existing ones for closed quantum systems  [40] and classical ones [18, 19, 20, 21].
@@
The superscripts highlight the sites on which the operators act. The first sum contains (possibly inhomogeneous) single-site terms $h_{k}$, while the second one describes long-range interactions, encoded in the (possibly inhomogeneous) two-body operators $V_{jk}$. The interaction strength decays algebraically with the distance, through the interaction-range exponent $\alpha$ [cf. Fig. 1(a)]. For $\alpha\leq d$, the double sum in Eq. (1) becomes super-extensive in $N$, leading to unphysical behavior in the thermodynamic limit $N\to\infty$ [26]. To prevent this, we include the so-called Kac factor $c_{\alpha}^{N}=\sum_{\textbf{r}\in[0,L-1]^{d},\textbf{r}\neq 0}\|\textbf{r}\|_{\Gamma}^{-\alpha}$ [see Eq. (1)], which restores the extensivity of the Hamiltonian [38]. For large $N$, the Kac factor scales as [26]
@@
$\gamma_{jk}^{\mu}=\kappa_{j}^{\mu}\delta_{jk}+\frac{w_{jk}^{\mu}}{2c_{\alpha}^{N}[\delta_{jk}+\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.$ (7)
@@
$\mathcal{L}_{N}=\sum_{k\in\Gamma}\mathcal{L}_{k}^{(k)}+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Gamma}\frac{\mathcal{L}_{jk}^{(jk)}}{[\delta_{jk}+\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.$ (8)
The second term features an unconstrained double sum, so that the factor $1/2$ avoids double counting. We assume $h_{k}$, $V_{jk}$, $\gamma^{\mu}$ and $L_{\mu}^{(k)}$ to be bounded.
@@
$\begin{split}\dot{\varrho}_{N,t}^{(\Lambda)}&=\sum_{k\in\Lambda}\mathcal{L}_{k}^{(k)}[\varrho_{N,t}^{(\Lambda)}]+\frac{1}{c_{\alpha}^{N}}\sum_{j\in\Lambda,k\in\Lambda^{\prime}}\mathcal{C}_{jk}^{N}[{\varrho}_{N,t}^{(\Lambda\cup\{k\})}]\\ &+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}[{\varrho}_{N,t}^{(\Lambda)}]}{[\delta_{jk}+\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,.\end{split}$ (9)
@@
for $j\in\Lambda$, $k\in\Lambda^{\prime}$. The relation in Eq. (9) defines a hierarchy of differential equations, manifesting in ${\varrho}_{N,t}^{(\Lambda)}$ depending on the reduced states ${\varrho}_{N,t}^{(\Lambda\cup\{k\})}$ involving the particles in $\Lambda$ plus a particle in $\Lambda^{\prime}$ [25, 28, 52, 51, 27].
@@
Our theorem shows that, for large $N$, reduced states of the many-body quantum system remain in product form, if they initially are, and are fully determined by the single-particle states $\tilde{\varrho}_{N,t}^{(j)}$. The latter evolve through a dissipative Hartree-like [25, 34, 51] dynamics [cf. Eq. (13)]. Contrary to a standard (permutation-invariant) mean-field scenario, the generic situation considered so far does not allow for an immediate statement on the convergence, with $N$, of $\tilde{\varrho}_{N,t}^{(j)}$ to a well-defined limit $\tilde{\varrho}_{\infty,t}^{(j)}$. This is due to the allowed inhomogeneous parameters, or disorder, in the dynamical generator [cf. Eq. (1) and Eq. (7)]. Moreover, it is intrinsically related to the fact that long-range interactions are not permutation invariant. As a consequence, the convergence of reduced single-particle states is strictly tied to the convergence of the average over collisions in Eq. (13) (see Proposition 1 below), which is not obvious as in permutation-invariant cases. Even in the most general case, our theorem however permits the efficient analysis of the system as well as of convergence of the reduced single-particle states, e.g., via numerical integration of Eq. (13) for increasing $N$.
@@
We take a spin-$1/2$ chain ($d=1$) with open boundaries, as illustrated in Fig. 2(a). The unitary part of dynamics is governed by a Hamiltonian like the one in Eq. (1), with $h_{k}=0$ and fixing $V_{12}^{(jk)}=\Omega\sigma^{(j)}_{x}\sigma^{(k)}_{x}$. The dissipative contribution is local, with
@@
with $\phi(\textbf{x})=\langle\sigma_{x}\rangle\Omega\theta(\alpha)\int_{0}^{1}{\rm d}\textbf{y}\,|\textbf{x}-\textbf{y}|^{-\alpha}$. Such an equation is benchmarked in Fig. 3(b).
The nonlinear and non-interacting character of the dynamics governed by Eq. (17) provides simple, yet rigorous, insights on the emergence of the long-lived quasi-stationary states observed in Refs. [26, 12]. There, a unitary quench from a state with $\langle\sigma_{x}\rangle=0$ was considered. Our results imply that, in the thermodynamic limit, the system must obey Eq. (17), with $\kappa=0$, so that an initial state with $\langle\sigma_{x}\rangle=0$ is a fixed point of the emergent dynamics. Therefore, for finite systems, the quantum many-body state must necessarily become longer and longer lived upon increasing the number of particles. [Fig. 3(b) displays dynamics since we consider a state with $\langle\sigma_{x}\rangle\neq 0$.] This example also illustrates the non-commuting character of the limits $N\to\infty$ and $t\to\infty$.
Discussion.— While exact results exist for Hamiltonian systems with collective [25, 40] and power-law interactions [56, 26, 27, 12, 40] as well as for open quantum systems with collective interactions [28, 29, 31, 32, 33, 34, 35, 37], our theorem goes beyond these cases. Furthermore, our work differs significantly from the ones mentioned above since Theorem 1, which takes inspiration from Refs. [25, 28, 51], is not restricted to homogeneous interactions, but covers the cases of disordered single-site and interacting terms as well as of generic inhomogeneous dynamics. Importantly, it allows for the exploration of strong long-range interacting dynamics with a linear complexity in the system size [cf. Eq. (13)]. Within this formalism, it is possible to describe nonequilibrium phase transitions and spontaneous symmetry breaking (see discussion in Ref. [50]). Our proposition demonstrates the impact of boundary conditions on the emergent dynamics. For periodic boundaries and translation-invariant dynamical generators, the system becomes permutation invariant in the thermodynamic limit. For open boundaries, it is instead described by an inhomogeneous mean-field theory (see, e.g., Refs. [57, 58, 59, 60] for inhomogeneous mean-field equations in related contexts), where contrary to recent results [61] the inhomogeneity emerges from the dynamics itself and is not a remnant of the chosen initial state. While Eqs. (13-15) may look abstract at first, they simply entail that the relevant equations of motion can be found by considering a product-state ansatz. We stress that our results are valid for reduced states and do not preclude the possibility of highly entangled full many-body states [62, 63] nor of measurement-induced phase transitions [64]. We also note that they can be exploited to compute the expectation value of quasi-local operators and not only of strictly local ones [65, 50].
@@
The code and the data for the dissipative quench described in the manuscript are available on Zenodo [67].
Acknowledgments.— We are grateful to Alvise Bastianello and Stefan Teufel for useful comments. We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Unit FOR 5413/1, Grant No. 465199066, through the Research Unit FOR 5522/1, Grant No. 499180199. This project has also received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). This work was funded within the QuantERA II Programme (project CoQuaDis, DFG Grant No. 532763411) that has received funding from the EU H2020 research and innovation programme under GA No. 101017733. FC is indebted to the Baden-Württemberg Stiftung for the financial support of this research project by the Eliteprogramme for Postdocs.
@@
Robert Mattes,1 Igor Lesanovsky,1,2 and Federico Carollo3
@@
3Centre for Fluid and Complex Systems, Coventry University, Coventry, CV1 2TT, United Kingdom
@@
$\mathcal{L}^{(\Lambda)}=\sum_{k\in\Lambda}\mathcal{L}^{(k)}_{k}+\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}}{[\delta_{jk}+\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\,,$ (S3)
@@
We note that the last sum can be rewritten as $\sum_{\Delta\textbf{r}\neq 0}||\Delta\textbf{r}||_{\Gamma}^{-\alpha}$, with $\Delta\textbf{r}=\textbf{r}_{j_{i}}-\textbf{r}_{k_{i}}$, which is similar to the definition of the Kac factor but with different summation indexes. In each dimension the summation indeed splits into the two parts $\textbf{r}_{j_{i}}^{\mu}-1,\dots,0$ and $-1,\dots,\textbf{r}_{j_{i}}^{\mu}-L$ and by adding terms to both parts in every dimension one finally recovers $2^{d}$ times the Kac factor. Moreover, each sum over $j_{i}$ is bounded by $n+i-1$, where $n=|\Lambda|$ is the extension of the subset $\Lambda$. All these considerations give
@@
$\left\|e^{s\mathcal{L}^{(\Lambda)}}\circ\left[\mathcal{L}^{(\Lambda)}-\hat{\mathcal{L}}^{(\Lambda)}\right]\circ e^{(t-s)\hat{\mathcal{L}}^{(\Lambda)}}[Q^{(\Lambda)}]\right\|_{\rm tr}$ $\leq\left\|\frac{1}{2c_{\alpha}^{N}}\sum_{j,k\in\Lambda}\frac{\mathcal{L}_{jk}^{(jk)}\circ e^{(t-s)\hat{\mathcal{L}}^{(\Lambda)}}[Q^{(\Lambda)}]}{[\delta_{jk}+\|\textbf{r}_{j}-\textbf{r}_{k}\|^{\alpha}_{\Gamma}]}\right\|_{\rm tr}\leq\left\|Q^{(\Lambda)}\right\|_{\rm tr}\frac{C|\Lambda|^{2}}{c_{\alpha}^{N}}\,.$
@@
$\|\mathcal{M}_{t}[\cdot]\|_{\mathrm{tr}}\leq\|\mathcal{L}_{1}\left[\cdot\right]\|_{\mathrm{tr}}+\int{\rm d}\textbf{z}\frac{1}{\|\textbf{x}-\textbf{z}\|_{\Gamma}^{\alpha}}\left\|\mathrm{tr}_{\textbf{z}}\circ\mathcal{L}_{12}\circ\left[\cdot\otimes\tilde{\varrho}_{\infty,t}^{(\textbf{z})}\right]\right\|_{\mathrm{tr}}\leq\left(E+\int{\rm d}\textbf{z}\frac{1}{\|\textbf{x}-\textbf{z}\|_{\Gamma}^{\alpha}}C\right)\|\cdot\|_{\mathrm{tr}}=M\|\cdot\|_{\mathrm{tr}}\,,$
@@
IV IV. Additional results for the long-range dissipative quench
@@
This formula shows that we can calculate the evolution of the reduced state $\varrho_{N,t}^{(\{j,k\})}$ as follows. First, we consider the interaction $\mathcal{U}_{jk}$ between the particles of interest. Then, we make both site $j$ and site $k$ sequentially interact with one particle in $\{j,k\}^{\prime}$ and tracing the latter out. After all particles in the remainder of the system have been considered, we can apply the dissipative contributions. This algorithm, schematically illustrated in Fig. (2)(b), allows us to investigate the many-body system with a complexity which is equivalent to that of a three-qubit system subject to order $N$ maps, which involve at most two qubits. In the numerical calculations, we have used the initial state $\varrho_{N}^{(j)}=\left(\begin{array}[]{cc}0.2&0.2-0.2i\\ 0.2+0.2i&0.8\end{array}\right)\,,$ for each site. In Fig. S1(a,b), we show results for the quantum mutual information between the two sites, $\mathcal{I}(j:k)=\mathcal{S}(\varrho_{N,t}^{(\{j\})})+\mathcal{S}(\varrho_{N,t}^{(\{k\})})-\mathcal{S}(\varrho_{N,t}^{(\{j,k\})})$, with $\mathcal{S}(\varrho)=-\mathrm{tr}\left(\varrho\log(\varrho)\right)$ being the von Neumann entropy. We see that the quantum mutual information decreases with the system size in the strong long-range regime $\alpha\leq 1$. Its behavior as a function of $N$ [shown in Fig. S1(b)] is in principle not the same shown in the inset of Fig. 3(b) by the correlations $C_{zz}$, due to the definition of the quantum mutual information through entropy functions.
V V. Extension to quasi-local operators
Beyond statements about expectation values of local observables, Theorem 1 immediately applies also to non-strictly local observables and correlation functions.
To understand this, we can indeed consider the so-called “quasi-local algebra” of a many-body system [65], which contains operators that are not strictly local but are still convergent in the norm topology. These quasi-local operators $A$ can be approximated, with arbitrary precision, by strictly local operators $A_{\epsilon}$. This implies that we can use our results to show the noninteracting character of the many-body state also on operators which are almost localized yet extended to the whole many-body system. This is because we can always find a suitable local operator $A_{\epsilon}$ such that $|\langle A_{\epsilon}\rangle-\langle A\rangle|<\epsilon\,,$ with $\epsilon>0$ small at pleasure, and given the fact that the expectation on local operators $\langle A_{\epsilon}\rangle$ shows a noninteracting character.
VI VI. Spontaneous symmetry breaking on the level of the emergent locally noninteracting dynamics
First of all, we note that in our approach, we first take the limit $N\to\infty$ which gives rise to the mean-field equations of motion for the systems. Starting from a dynamical (Lindblad) generator obeying a given symmetry, one also has that the mean-field equations obey the same symmetry of the generator. In general, the mean-field equations of motion have a unique stable solution which also obeys the same symmetry of the generator. Spontaneous symmetry breaking in these long-range interacting systems occurs when the mean-field equations feature two, or more, stable solutions which break the symmetry of the generator (see e.g. Ref. [30]). Each of these solutions is dynamically approached according to which initial state is chosen for the system.
Importantly, we note that mean-field dynamics are typically valid only when the initial state of the system has low correlations (in our case we considered initial product states). These states are usually not symmetric and they thus approach only one of the different stable solutions available in the symmetry-broken phase. The way to restore the symmetry in this setup is to consider the state given by the sum of different (product) mean-field solutions, associated with different initial states, such that the incoherent mixture state obeys the symmetry of the generator. This leads to a stationary state that is given by the sum of all stable (product) solutions and is thus symmetric. In the complementary approach, in which one is able to first take the limit $t\to\infty$, the stationary state of the system is symmetric since for any finite $N$ there is no symmetry breaking. Spontaneous symmetry breaking then solely occurs when taking the limit $N\to\infty$, manifesting through the emergence of an incoherent mixture state which, at the level of local observables, behaves as the incoherent sum over the (product) stable solutions of the mean-field equations.
To concretely understand the above mechanism, let us consider the simple mean-field equation of the magnetization $m$ in the classical Ising model. The dimensionless version of the equation is given by
$\dot{m}=-m+\tanh(\beta m)\,,$ (S24)
where $\beta$ is a dimensionless inverse temperature. The above equation is invariant for $m\to-m$, which constitutes the $Z_{2}$ symmetry of the model. For sufficiently small values of $\beta$ the system features a unique solution $m=0$. For $\beta$ large enough, we instead have that two possible solutions $m_{1}$, $m_{2}$ exist, with opposite magnetization $m_{1}=-M$ and $m_{2}=M$. In this symmetry-broken regime, one can see that initial states with negative magnetization will approach solution $m_{1}$, while those with positive magnetization will approach $m_{2}$. To construct a symmetric state, one can consider the sum of an initial (product) state with $m=-\bar{m}$ plus another initial (product) state with $m=\bar{m}$. The latter incoherent initial state converges to a stationary state given by the sum of the (product) state with magnetization $-M$ and the (product) state with magnetization $M$. Note that this is different from having a stationary state made by a single (product) state with magnetization equal to zero.